

********************************************************************
UARS WEEKLY STATUS REPORT - 23 Nov 2001
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues normal operations, taking solar data on all available
orbital tracking periods and stellar calibration observations whenever
possible. We continue to gather enough data to meet the SOLSTICE science
objectives during this period of reduced tape recorder operations.
***********
SUSIM
***********
SUSIM completed daily mid resolution and MgII solar scans
during periods of predicted UARS telemetry.
We continue work on the remaining software components that will allow
the processing of the SUSIM data from level 1 to level 3 on the SUSIM
RAC.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact
for UARS instrument operations to support re-engineering
of the UARS Mission Operations Concept to develop
capabilities for the Science Traceability Mission.
Daily processing of instrument Daily Activity Plans
continues with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 10 November 2001 (Orbit 55577, GMT Day 314) through
16 November 2001 (Orbit 55681, GMT Day 320).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 317/1248 55629 Power ON
ACRIM 320/1605 55677 Power OFF - leonids
The beta angle ranged between the angles of 65.2 degrees to 42.8 degrees
for this report period. The beta angle is now increasing towards a
minimum angle of 2 degrees on 26 November 2001 (DOY 330).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.8 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 1.0 to 1.3 degrees C. Additional
battery performance data for each battery on 09 November 2001 and 16
November 2001 are provided in Appendix D.
The clock error ranged from -12.5 to 10.4 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and
Out-Of-Limits occurrences are listed in Appendix B.
95 OUT-OF-LIMITS 3D 00
95 ALERTS 3D 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
95 FOT 3D 17
95 PEM 3D 1
95 HALO 3D 1
95 ACRIM 3D 1
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
95 FOT 3D 02
One (1) Anomaly Report remained OPEN at the end of this report period
(see Appendix E).
TRANSPONDER FREQUENCIES:
* Transponder A: 00/316 - 2287.496493
* Transponder B: 01/286 - 2287.498613
GROUND SYSTEM OPERATION
Two (2) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data Capture
Facility (DCF). This service should resume soon.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT and software developers located the cause the problems on the
new TAC system. A temporary patch was installed for testing. Testing
showed that the patch resolved the urgent DRs. The TAC developers are
now in the process of fixing the software permanently and will delivery
a new release sometime early next week. This should allow the new TACs
to go to ORR and subsequently released for operations.
MEETINGS
The FOT attended a UARS Eclipse Meeting on Wednesday, November 15, 2001.
The FOT attended a UARS OBSTF on Wednesday, November 15, 2001.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 26 November 2001 (DOY 330).
This maneuver will be a Reverse to Forward Flight maneuver (Southern
Hemisphere viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 16 November 2001
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
55615 316/1324 01-666 OBC Clk Adj fr 20.6 to 0.00 msec/day
55618 316/1829 01-667 NBTR NOOPS D-317
55622 317/0120 01-668 PEM TO DAY ONLY OPS
55626 317/0728 01-667 NBTR NBTR PBK & RTS-58
55627 317/0809 01-667 NBTR/OBC NOOPS D-317
55629 317/1248 01-669 HALOE Power ON
55633 317/1827 01-670 NBTR NOOPS D-318
55640 318/0544 01-670 NBTR NBTR PLAYBACK
55642 318/0809 01-670 NBTR/OBC NOOPS D-318
55645 318/1325 01-670 NBTR NBTR PLAYBACK
55648 318/1826 01-671 NBTR/OBC NOOPS D-319
55649 318/2005 01-672 NBTR NBTR TO REC MODE
55654 319/0402 01-673 OBC/CLK Clk Adj 0.00 to 20.6 msec/day
55655 319-0544 01-671 NBTR NBTR PBK & RTS-58
55657 319/0808 01-671 NBTR/OBC NOOPS D-319
55663 319/1747 01-671 NBTR NOOPS D-319
55663 319/1825 01-674 NBTR NOOPS D-320
55670 320/0627 01-674 NBTR NOOPS D-320
55675 320/1425 01-675 OBC Clk Adj fr 20.6 to 0.00 msec/day
55677 320/1605 01-676 ACRIM Power OFF - leonids20
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 16 November 2001
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
_________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 16 November 2001
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
-None this report period.
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 16 November 2001
09 November 2001 - GMT Day 313 Beta 3D -68.0 deg, SA Pos 3D 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 +16.8/+00.0 -3.68/-4.62 +0.0/+0.0 13.0 0.000 0.0
2 +16.8/-44.8 +5.15/+3.87 +26.8/-15.2 25.1 1.030 20.5
3 +22.4/-67.2 +4.51/+2.92 +26.4/-15.2 25.1 1.020 20.2
16 November 2001 - GMT Day 320 Beta 3D 42.8 deg, SA Pos 3D 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- ---------- ---------- ----- ------- ------
1 +728.0/+728.0 -3.77/-4.38 +0.0/ -0.0 10.7 0.000 0.0
2 +00.0/ -72.8 +5.47/+4.48 +28.4/-14.4 25.8 1.050 20.4
3 +22.4/ -22.4 +4.48/+3.17 +27.2/-14.4 25.8 1.040 20.0
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---- ---------
-no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 16 November 2001
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-072 FOT 55654 319/0449 24106 Dropout - NO CARRIER
01-073 FOT 55678 320/1823 N/A Ops Ctr Late AOS
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-072 FOT 55654 319/0449 24106 Dropout - NO CARRIER
01-073 FOT 55678 320/1823 N/A Ops Ctr Late AOS
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-69 FOT see list Nov 01 14170 Generic Late Acquisition
SEVNTFW01/46